One kilogram of fresh air is compressed isothermally and reversibly from
1.0 bar and 30 oC to 5.0 bar. Calculate the work done w on the air and
the heat flow q to or from the air.
1
Expert's answer
2020-12-28T04:42:28-0500

"W=\frac{m}{M}RT\ln{\frac{p_1}{p_2}}=\frac{1}{0.029}\cdot 8.31\cdot 303\cdot\ln{\frac{100000}{500000}}=-140(kJ)"


"Q=\Delta U+W=0+W=-140(kJ)"